OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -9) and a directrix y = 9.

OpenStudy (anonymous):

these are the answer choices; y = -1/9 x2 y2 = -36x y = -1/36 x2 y2 = -9x

OpenStudy (anonymous):

so would it be the y^2 = -36x one?

OpenStudy (anonymous):

lol thank you so much! would u help me with one more.. these are my practice questions before my assignment lol

OpenStudy (anonymous):

post it

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3.

OpenStudy (anonymous):

here are the answer choices; y = 1/12 x2 -12y = x2 x = 1/12 y2 y2 = 6x

OpenStudy (anonymous):

so would it be the -12y = x^2?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

wait? what ? thats not one of the answer choices

OpenStudy (anonymous):

ahh!! i seee! thank you!!